OpenStudy (anonymous):

if you pay rent $2000 incl 15% Vat how much is your rent without Vat?

OpenStudy (anonymous):

Lets say your rent without vat is x. Then, \[(x+0.15*x)=1.15*x=2000\]

OpenStudy (anonymous):

thanx